OpenStudy (anonymous):

Which of the following shows the factors of 6x2 - 3x - 18?

OpenStudy (cggurumanjunath):

3(2x^2-x-6)=3(2x^2-4x+3x-6)=3(2x(x-2)+3(x-2))=3(2x+3)(x-2)